OpenStudy (anonymous):

i need HELP....If f(x) = x 2 + 1 and g(x) = 3x + 1, find 2·f(4).

OpenStudy (austinl):

We have, \(f(x) = x 2 + 1\) We want, \(2\times f(4)\) That means that you plug in for for all instances of x in f(x), and then multiply that by 2.

OpenStudy (anonymous):

I don't know how u mean...

OpenStudy (austinl):

\(f(4)=(4)^2+1=17\rightarrow 17\times2=?\)

OpenStudy (anonymous):

thanks that really helped...:)

OpenStudy (austinl):

You are welcome!